40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1495     Entry time: Sat Aug 30 19:10:11 2014
Author: Evan 
Type: DailyProgress 
Category: NoiseBudget 
Subject: AlGaAs python noise budget: seismic coupling added 

Same data and same isolation model as for the silica/tantala noise budget. Since we have new table legs, we should retake this data (and make a spectral histogram).

The resonance frequencies of the stack are given as 10 Hz and 35 Hz in the noise budget. Are these for the old stack? I recall that with the new stack we measured resonances at 3, 7, and 10 Hz.

Also I want to double check the sequence of interpolation steps we've used on the silica/tantala noise budget. There are some seismic peaks and silica/tantala beat peaks that almost (but don't quite) match up in frequency, and I wonder whether this is an artifact of the interpolation.

Attachment 1: noiseBudget.pdf  192 kB  | Hide | Hide all
noiseBudget.pdf
ELOG V3.1.3-